OpenStudy (anonymous):

Find the complex roots in the following: f(x) =(x-1)(x^2-3)(x^2+3)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

complex roots are the ones you get from setting \[x^2+3=0\] you solve in two steps only do you know how?

OpenStudy (anonymous):

Not really. Can you tell me how to find it?

OpenStudy (misty1212):

sure a a) subtract \(3\) from both sides, get \[x^2=-3\]

OpenStudy (misty1212):

b) then take the square root of both sides \[x=\pm\sqrt{-3}\]

OpenStudy (xapproachesinfinity):

x^2+3=0 subract 3 from both sides x^2=-3 take square root of both sides x=+ or _ root(-3)

OpenStudy (xapproachesinfinity):

eh misty got this :)

OpenStudy (misty1212):

because the number inside the radical is negative , it is not a real number you can write it as \[x=\pm\sqrt3 i\] if you like

OpenStudy (anonymous):

So the complex roots are 1, \[\pm \sqrt{3}\]?